OpenStudy (anonymous):

In right triangle ABC, ∠A and ∠B are complementary angles and sin A = 89. What is cos B? 817−−√17 89 17−−√9 17−−√8

OpenStudy (anonymous):

8√17/17 8/9 √17/9 √17/8

geerky42 (geerky42):

\(\sin\theta = \dfrac{\text{Opp}}{\text{Hyp}}\) \(\cos\theta = \dfrac{\text{Adj}}{\text{Hyp}}\)